### v1

# Simple model for the kinetics of stimuli-responsive gels with porous structures

Abstract
Stimuli-responsive gels are the gels that vary the volume depending on environmental conditions. It has been reported that the stimuli-responsive gel with an inhomogeneous structure exhibits faster volume change than the gel without it. It is understood as a difference in the transfer dynamics of the solvent, though, there are few models for discussing the effect of inhomogeneity explicitly. In this paper, we propose a simple model for the kinetics of volume change by introducing inhomogeneity as the probability distribution for a random variable that characterizes the structure. Under this framework, we demonstrate that inhomogeneity actually increases the rate of volume change. Furthermore, through analysis of a simplified example, we show that some features that are often seen in experiments can be reproduced by the model. This model provides a simple equation for the time series of volume change that can be used to characterize the experimental data based on a simple physical picture of the phenomenon.
I Introduction
Stimuli-responsive gels are the gels that vary their volume depending on environmental condition such as temperature, pH, irradiation, etck. Ahn et al. (2008). In recent years, they have attracted considerable attention in engineering and biomedical fields as smart materials that adapt to the environmentHines et al. (2017); Lavrador et al. (2021). In many applications of the stimuli-responsive gels, there is demand of increasing ratio and rate of volume change. The basic mechanism of the volume change is explained by the Flory-Huggins theory, which is a statistical model that gives the osmotic pressure of gels by calculating the free energy of the gel state while considering the entropy and enthalpy for mixing polymer and solventDoi (2013). The volume of gel is determined to balance the osmotic pressure and elastic stress of polymer network. In the case of the stimuli-responsive gel, the environmental condition changes the interaction between the polymer and solvent, and it changes the osmotic pressure in the gel. Then, the gel absorbs or expels the solvent, which results in a change in volume.
In this mechanism, the rate of volume change is limited by either the transfer rate of stimulus (like heat) or the rate of polymer network’s motion. In most cases, the motion of the polymer network is slower than the transfer of stimulus; thus, it becomes a rate-limiting step of volume changeGehrke (1993). The motion of a polymer network is well described by the Tanaka-Fillmore theory, in which the motion of a polymer network follows a diffusion processTanaka et al. (1973); Li and Tanaka (1990); Doi (2009). The rate of diffusion is characterized by the diffusion coefficient $D$, which equals to the ratio of the elastic modulus of the polymer network to the friction coefficient between the network and solvent. Many efforts are devoted to enhance $D$ through control of the synthesis method and chemical conditions, though, recent studies reveal another approach is very effective in increasing the rate of volume change; introducing inhomogeneity. Experimental studies revealed that when a responsive gel is made with a microporous open-celled structure, the solvent can be absorbed or expelled by the convection process, which makes the swelling/shrinking kinetics much faster than that by the diffusion process of polymer networkKabra et al. (1992); Takata et al. (2002); Kato and Gehrke (2004); Gotoh et al. (2004); Maeda et al. (2008); Murase et al. (2009); Seiffert (2012). It is also reported that inhomogeneous structure drastically changes static properties of gels like equilibrium swelling ratio and elasticityF. Di Lorenzo (2015); Lopez et al. (2019). From theoretical perspective, while great development of numerical simulation method enables to explore precise dynamics in the internal of responsive gelsNikolov et al. (2018); Camerin et al. (2018), theoretical understanding based on a simple model that captures essential features seems to be less progressed. To understand the role of inhomogeneity in the kinetics, we need a model that explicitly treats inhomogeneity. In this study, we propose a simple model for the kinetics of responsive gels that introduces inhomogeneity as a probability distribution. Under the framework, we first discuss the role of the inhomogeneous structure and show that the inhomogeneity actually makes the kinetics faster. Furthermore, we show that some features that are often seen in experiments can be reproduced by the model.
II Model
We focus on the uniform deformation of gels. Then, the volume change in infinitesimal time variation is equal to the flow rate through the surface. Denoting the volume of gel as $V$, the flow rate of solvent into the gel as $Q$ and, time as $t$,
$\frac{dV}{dt}=Q.$ (1)
Our strategy is to model the flow of solvent in porous gels by the Hagen-Poiseuille flow and introduce an inhomogeneous structure as the distribution of the flow channel. In the absorbing and expelling process of the solvent in actual gels, the solvent flows through tortuous channels. Here, we approximate the channels as a bundle of capillaries with different radii. Denoting the radius of capillary $j$ as $\lambda_{j}$ and flow rate of that as $q(\lambda_{j})$, which is positive for the direction from external to internal of the gel, the total flow rate into the gel is
$Q=\sum_{j}q(\lambda_{j}).$ (2)
For $q(\lambda_{j})$, we assume the Hagen-Poiseuille law, which is plausible for flow in cylindrical pipes at low Reynolds numbers, as follows:
$q(\lambda_{j})=\gamma\frac{\lambda_{j}^{4}}{l}\Delta p,$ (3)
where $l$ is the length of the channel, $\Delta p$ is the pressure difference between the external and internal of the gel, and $\gamma$ is a constant. The approximation for $Q$ is similar to that used for modeling flow in solid porous media, which is found to be valid for the casesWu and Yu (2007); Yu (2008). $\Delta p$ is computed from various factors reflecting each mechanism of stimuli-responsiveness and is given for individual cases. Although $l$ and $\Delta p$ can be different for each channel, we assume them to be identical for simplicity here. Note that $\gamma,\lambda_{j}$ and $l$ are positive from their definition, though, $\Delta p$ can be positive and negative. The sing of $\Delta p$ determines the direction of flow and so we can apply this framework to both swelling and shrinking process.
For further simplification, we assume that the shape of the flow channel is scaled by the total volume of channels (which equals to the volume of solvent) as
$\lambda_{j}=A_{j}V_{s}^{\frac{1}{3}},\quad l=BV_{s}^{\frac{1}{3}},$ (4)
where $V_{s}$ is the volume of solvent and $B$ and $A_{j}$ are positive constants. $A_{j}$ are random numbers and we denote the probability density function as $p(A)$. $p(A)$ represents the inhomogeneity of the structure. Denoting the volume of the polymer as $V_{p}$,
$V$ $=V_{p}+V_{s}$
$=V_{p}+\sum_{j}\pi\lambda_{j}^{2}l$
$=V_{p}+\pi BV\sum_{j}A_{j}^{2}.$ (5)
When the number of $j$ is sufficiently large, the summation is replaced by an integral. Then, from Eq. (5) we obtain
[…]
This becomes the normalization condition for $p(A)$ to assure that $V,V_{p}$ and $B$, i.e., the whole volume, volume of polymer and length of polymer, are fixed. (Note that when $V$ and $V_{p}$ are given, $V_{s}=V-V_{p}$ is fixed and so fixing $B$ is equivalent to fixing $l$.) Substituting Eqs. (2), (3), (4) into Eq. (1) and replacing te summation with an integral, we obtain
[…]
When the volume fraction of polymer in the gel is sufficiently small, (which is often valid for responsive gels,) $V_{p}$ becomes negligible compared with $V$. Then, we obtain
[…]
This becomes the differential equation for $V$ when $\Delta p$ and $f(A)$ are given.
Now, we consider a case that $\Delta p$ is composed of osmotic pressure by chemical concentration and resilience force by elasticity of structure. For simplicity, We assume $\Delta p$ only depends on $V$ as
$\Delta p(V)=\frac{\alpha}{V}-\beta\left(\frac{V-V_{n}}{V_{n}}\right),$ (9)
where $\alpha,\beta$ and $V_{n}$ are constants that represent the strength of osmotic pressure, strength of resilience force, and natural volume of the gel, respectively. To consider physically meaningful cases, we assume $\beta>0$ and $V_{n}>0$. Substituting Eq. (9) into Eq. (8), we obtain
$\frac{dV}{dt}=\gamma^{\prime}\left\{\alpha-\beta\frac{V}{V_{n}}\left(V-V_{n}\right)\right\},$ (10)
where
[…]
This equation is equivalent to the logistic equation. Although there are four parameters in Eq. (10), since $\alpha,\beta$ and $\gamma^{\prime}$ attribute to the equation only in the form of $\alpha$ and $\gamma^{\prime}\beta$, the number of relevant parameters is three. Figure 1 shows the schematics of the phase space of Eq. (10). $V_{e}$ indicates the equilibrium volume corresponding to a fixed point of the dynamics. Requiring the existence of a stable fixed point, $\alpha>-\beta V_{n}/4$ must be assumed. When $\alpha<0$, there is a solution that $V\to-\infty$. This is physically nonsense solution due to the simplicity of the assumption for $\Delta p$. Thus, we ignore such an initial condition in the following discussion.
Figure 1: Phase space of the differential equation (10). Red arrows show the direction of vector field.
III Analysis
With the initial condition ($V(0)=V_{0}$), we can integrate the differential equation (10) as
$V(t)=\frac{-a}{ce^{t/\tau}+1}+V_{e},$ (12)
where
$\begin{split}a=V_{n}\left(1+\frac{4\alpha}{\beta V_{n}}\right)^{\frac{1}{2}},\quad\tau=\frac{1}{\gamma^{\prime}\beta}\left(1+\frac{4\alpha}{\beta V_{n}}\right)^{-\frac{1}{2}}\\ V_{e}=V_{n}\left\{\frac{1}{2}+\frac{1}{2}\left(1+\frac{4\alpha}{\beta V_{n}}\right)^{\frac{1}{2}}\right\},\quad c=\frac{V_{e}+V_{0}-V_{n}}{V_{e}-V_{0}}.\end{split}$ (13)
The four parameters described by Eqs. (13) correspond to the set of the three relevant parameters in Eq. (10) and the initial value of $V$. $V_{e}$ suffices $\lim_{t\rightarrow\infty}V(t)=V_{e}$ as well as $\Delta p(V_{e})=0$. When $\alpha/\beta V_{n}$ is small, $V_{e}\simeq V_{n}+\alpha/\beta$, which means that the equilibrium volume is the sum of the natural volume and the ratio of the strength of the osmotic pressure and the resilience force. $\tau$ characterizes the relaxation time of the exponential dependency, which decreases when $\gamma^{\prime}$ and $\alpha$ becomes larger.
Figure 2 shows three typical behaviors of the model. Parameters are set as $\alpha=0.1,\beta=10,\gamma^{\prime}=1,V_{n}=1$, and the case for three initial conditions, $V_{0}=0.1,0.6,1.6$, are depicted. From Eq. (13), $V_{e}\approx 1.1$ and $\tau\approx 0.1$ for these cases. Each curve represents the typical behavior of large swelling, small swelling, and shrinking. When $V_{0}$ is sufficiently small, $V(t)$ has an inflection point at which the sign of curvature changes. When $V_{0}$ is not so small, $V(t)$ does not have an inflection point. From Fig. 1, it is found that the inflection point is $V=V_{n}/2$, so the condition to have an inflection point is $V_{0}/V_{n}<1/2$. When $c\gg 1$, the solution (12) becomes
$V(t)\simeq-\frac{a}{c}e^{-t/\tau}+V_{e},$ (14)
which indicates simple exponential relaxation.
Figure 2: Typical time evolution of volume in the model. ($\alpha=0.1,\beta=10,\gamma^{\prime}=1,V_{n}=1$ and $V_{0}=0.1,0.6,1,6$.)
We now discuss the effect of inhomogeneity on the rate of volume change. In our model, inhomogeneity is introduced as the randomness of $A_{j}$, which characterizes the radius of cylindrical pores. Consider the case where the probability density function is a Gaussian distribution:
$f(A)=\frac{C}{\sqrt{2\pi\sigma^{2}}}\exp{\left(-\frac{(A-\mu)^{2}}{2\sigma^{2}}\right)},$ (15)
where $\mu,\sigma$ and $C$ are constants that represent the mean, standard deviation, and normalization constant, respectively. To ensure $A_{j}>0$, we assume $\mu\gg\sigma$ here. Now, we compare the cases in which all physical parameters, $V,V_{p},l,\alpha,\beta,\gamma,V_{n}$ and $\mu$, are the same except $\sigma$. When $V,V_{p}$ and $l$ are fixed, we can compute $C$ by Eq. (6) as
$C=\frac{1}{\pi B}\left(1-\frac{V_{p}}{V}\right)\frac{1}{\mu^{2}+\sigma^{2}}.$ (16)
Then, $\gamma^{\prime}$ is computed from Eq. (11) as
$\gamma^{\prime}=\frac{\gamma}{\pi B^{2}}\left(1-\frac{V_{p}}{V}\right)\frac{1}{\mu^{2}+\sigma^{2}}(\mu^{4}+6\mu^{2}\sigma^{2}+3\sigma^{4}).$ (17)
It is found from Eq. (10) that the rate of volume change is proportional to $\gamma^{\prime}$. Bacause the variation of $\sigma$ only affects $\gamma^{\prime}$ in the right-hand side of Eq. (10), one can discuss the effect of $\sigma$ through comparing the value of $\gamma^{\prime}$:
$\gamma^{\prime}(\sigma\neq 0)-\gamma^{\prime}(\sigma=0)=\frac{\gamma}{\pi B^{2}}\left(1-\frac{V_{p}}{V}\right)\frac{5\mu^{2}\sigma^{2}+3\sigma^{4}}{\mu^{2}+\sigma^{2}}\geq 0.$ (18)
This indicates that the rate of volume change is larger for the inhomogeneous case ($\sigma\neq 0$) than for the homogeneous case ($\sigma=0$).
IV Discussion
Although Eq. (12) is derived by assuming a very simple description for $\Delta p$ as Eq. (9), it would be useful for characterizing the time series of volume change obtained experimentally. It is often reported that the time series of volume change, both swelling and shrinking process, is well fitted by exponential function Li and Tanaka (1990); Suzuki and Hara (2001). Equation (14), which is derived from Eq. (12) with $c\gg 1$, is consistent with the behaviors. From Eqs. (13), $c\gg 1$ is equivalent to $V_{0}/V_{n}\gg 1$. This suggests that the exponential relaxation of volume is observed when the elastic resilience force sufficiently affects the kinetics. Although it should be noted that the Tanaka-Fillmore theory can also reproduce such exponential relaxation Li and Tanaka (1990).
When $V_{0}$ is sufficiently smaller than $V_{n}$, there emerges an inflection point as discussed. This reflects the balance between the hydrodynamic and elastic effect: in the swelling process, an increase in volume reduces the hydrodynamic resistance of capillaries, which enhances the flow rate, whereas approaching $V$ to $V_{e}$ suppresses the driving force, $\Delta p$. In the model (Eq. (10)), the dominant effect changes at $V=V_{n}/2$. Such transition of the dominant effect would be universal even when $\Delta p$ is a more realistic complex function. Previous experimental studies of responsive gels have often reported that the time series of volume change exhibits non-single exponential behaviors such as a sigmoid curveKabra et al. (1992); Kaneko et al. (1995). These behaviors are explained through various ideas such as formation of skin layer, domain structure in gel, etc., and they seem to be reasonable for each case. However, here we suggest a possibility that such behavior can be observed in some cases without considering the special dynamics in the gel.
V Conclusion
In this paper, we propose a simple model for the volume change of stimuli-responsive gels with a porous structure. We approximated the structure using a bundle of capillaries with different radii and modeled the rate of volume change to be equal to the hydrodynamic flow rate. In the framework, the driving force is the pressure difference, $\Delta p$, between the internal and external of the gel, which should be given for individual cases. For example, we discussed the simplified case when $\Delta p$ consists of the osmotic pressure and elastic resilience force. From the analysis, we found that the inhomogeneity of the structure increases the rate of volume change. Furthermore, we found that depending on the initial volume, the time series of the volume exhibits the exponential relaxation or sigmoidal behavior. We expect that the model will explain some of the experimental results on the kinetics of stimuli-responsive gels.